OpenStudy (anonymous):

interval [0,2pi) sinx^2 - 6cosx+6=0

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

in the future you should write "sin squared" as sin^2(x) or (sinx)^2

myininaya (myininaya):

(sin^2x)=1-cos^2x 1-cos^2x-6cosx+6=0 -cos^2x-6cosx+7=0 cos^2x+6cosx-7=0 let u=cosx u^2+6u-7=0 (u-1)(u+7)=0 cosx=1 cosx=-7

myininaya (myininaya):

so solve both of those equtions

myininaya (myininaya):

if cosx=1 then x=0 cosx can't be -7 so x=0 is only solution

myininaya (myininaya):

in the interval [0,2pi)

OpenStudy (anonymous):

thankyou

myininaya (myininaya):

if you like my answer give me a medal :)